# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v1

@@ Abstract
The rapid advances of multi-modal agents built on large foundation models have largely overlooked their potential for language-based communication between agents in collaborative tasks. This oversight presents a critical gap in understanding their effectiveness in real-world deployments, particularly when communicating with humans. Existing agentic benchmarks fail to address key aspects of inter-agent communication and collaboration, particularly in scenarios where agents have unequal access to information and must work together to achieve tasks beyond the scope of individual capabilities. To fill this gap, we introduce a novel benchmark designed to evaluate the collaborative performance of multimodal multi-agent systems through language communication. Our benchmark features a variety of scenarios, providing a comprehensive evaluation across four key categories of agentic capability in a communicative collaboration setting. By testing both agent-agent and agent-human collaborations using open-source and closed-source models, our findings reveal surprising weaknesses in state-of-the-art models, including proprietary models like GPT-4o. These models struggle to outperform even a simple random agent baseline in agent-agent collaboration and only surpass the random baseline when a human is involved. [^1] We will release our benchmark and evaluation code upon acceptance.
@@ 1 Introduction
The field of multimodal agents is experiencing rapid growth  (Xu et al., 2024; Xie et al., 2024; Cao et al., 2024), with research efforts expanding at an unprecedented pace. However, amidst this growth, a critical gap in research has emerged: the lack of focus on collaborative work (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024) among multiple multimodal agents. Synergistic operation of such agents is a highly promising but largely unexplored domain. Language agents can collaboratively finish complex tasks such as software development by assuming functional roles such as system designer, function generator, etc  (Qian et al., 2023; Du et al., 2024)). Current research on multimodal agents (Xu et al., 2024; Xie et al., 2024; Cao et al., 2024) has predominantly focused on individual agent capabilities, neglecting the potential for inter-agent collaboration. This limitation is further compounded by existing benchmarks such as VisualWebArena (Koh et al., 2024) and MME-RealWorld (Zhang et al., 2024), which fail to adequately assess collaborative performance between agents. As a result, our ability to evaluate and improve multi-agent systems remains constrained, hindering progress in this crucial area.
@@
Several critical questions emerge in the context of multimodal agent collaboration. How can different agents effectively communicate multimodal information through language when they have varying levels of access to information? In scenarios where different agents possess diverse task-specific capabilities, how can they collaborate to accomplish objectives that are beyond the scope of any individual agent? These research settings remain largely uncharted and present significant challenges. Furthermore, the ability of agents to handle incomplete information is of paramount importance, particularly when working with sensitive data  (Li et al., 2024) (i.e. Agent application in healthcare where privacy concerns are critical  (Tang et al., 2023)). The exploration of these questions is crucial for advancing the field of multimodal agent collaboration. By addressing these challenges, we can expand the applicability of multimodal agents in real-world scenarios (Zhang et al., 2024), particularly those involving sensitive or restricted information.
@@
Figure 1: Overview of the interaction between the Solver and Expert agents in our benchmark. The game manager presents the Solver agent with a puzzle, where the Solver can choose to interact by clicking or requesting advice from the Expert agent. The Solver is shown an image of the puzzle (a KeyPad puzzle in this instance) and makes decisions based on the possible actions. The Expert, informed by instruction manuals, provides guidance based on the Solver’s descriptions, such as advising which buttons to press. The interaction between the Solver and Expert is captured in the dialogue, reflecting the cooperation necessary to complete the task. Both agents use self-reflection on their choices by being prompted with the conversation history as it progresses.
@@
Motivated by these aforementioned issues, we propose a novel benchmark for evaluating collaborative multi-modal multi-agent frameworks to address critical gaps in current approaches (see  Figure 1). Our evaluation setting simulates a scenario where an in-house agent with direct access to sensitive data (i.e., the AI solver) collaborates with external expert agents (i.e., the AI expert) to analyze information without compromising privacy. This evaluation setting could revolutionize how we handle and extract insights from sensitive datasets across various domains.
@@
We assess multi-modal multi-agent systems using a series of carefully designed collaborative puzzle games. These scenarios typically involve two-player setups where agents have access to different, complementary information. (i.e., in a bomb defusal game, one agent possesses details about the bomb, while the other has access to a disarming manual). By employing such diverse and interactive scenarios, we aim to provide a thorough assessment of multi-modal multi-agent performance.
@@
Our benchmark includes 10 distinct, easily customizable puzzles with thousands of unique solutions. We tested two different settings (AI-AI and AI-Human) and evaluated several popular multimodal models, including closed-source models (GPT-4V, GPT-4O, and GPT-4o1) and open-source models (Qwen-VL (Bai et al., 2023), and InternVL(Chen et al., 2024)). Surprisingly, the GPT series does not outperform even a simple random baseline in the AI-AI setting, highlighting a potential growth area for future model development. Our contributions are as follows:
@@
  • Using COMMA, we carefully record conversations and performance metrics between state-of-the-art multimodal models such as QWenVL, InternVL, GPT-4o, GPT-4o1, etc (Section 4).
  • We categorize the agent capabilities tested in our model and common failure modes, providing insight into future research directions for improving inter-agent communication (Section 5).
@@ 2 Related Work
There are many emergent agent collaboration works  (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024; Ghafarollahi & Buehler, 2024; Li et al., 2023; Wu et al., 2023) among multiple language agents. Multi-agent systems arise mainly in two different scenarios: (1) role-playing different task executors (e.g., software development requiring different roles of agents, such as program manager, software architect, programmer  Du et al. (2024); Qian et al. (2023); Hong et al. (2024), scientific discovery simulation  Wu et al. (2023), and social simulation  Park et al. (2023); Gurcan (2024)); (2) communicating between agents with different pieces of information  Wu et al. (2023); Li et al. (2023) (e.g., consulting experts without sharing some sensitive or confidential data. In our case, the AI solver has some private multimodal data, and the AI expert has domain-specific knowledge or instructions).
@@
Instruction-based agent benchmarks evaluate an agent’s capability of following a human instructions to finish a task (e.g., navigating on a website, interacting with an operating system  Xu et al. (2024); Xie et al. (2024); Cao et al. (2024)). However, our benchmark focuses more on a communication-based evaluation where two clients engage in multi-turn conversations to solve a task collaboratively.
@@ 3.1 Design Principles of the Benchmark
Our benchmark is inspired by the cooperative gameplay scenario in Keep Talking and Nobody Explodes Games (2015). In this game, two players work together to defuse a bomb under time pressure. One player, the defuser, can see the bomb but lacks the instructions to disarm it. The other player, the expert, has access to the bomb’s manual but cannot see the bomb itself. The players must rely on effective communication to exchange information, navigate challenges, and defuse the bomb.
We generalize this dynamic for our benchmark, shifting the focus to solving vision-language puzzles in a communication-based agent framework. As multimodal agent systems gain momentum, our goal is to create a benchmark that rigorously evaluates their reasoning, communication, and collaborative abilities. The design of our benchmark revolves around the following core principles:
@@
A critical aspect of our benchmark is evaluating natural language communication between agents. Similar to how players in the original game exchange information verbally, agents in our framework must use language to share observations, clarify ambiguities, and reason about tasks. In order for the agents to succeed, they must display clarity, efficiency, and depth of communication, making it an essential factor in task completion.
Multi-agent collaboration:
In our benchmark, agents must work together, much like the two-player dynamic of Keep Talking and Nobody Explodes. The collaboration element ensures that tasks require mutual dependency, where each agent contributes unique information or capabilities that are critical to solving the puzzle. This principle highlights how well agents can cooperate and leverage each other’s strengths.
@@ 3.2 Categories of Agent Capability
Real-time Reaction (RT) Some puzzles challenge agents to process information rapidly and act with precise timing. This is a critical skill for embodied agents operating in dynamic, real-world environments where precise timing and quick reactions are vital.
@@ 3.3 Tasks
  • ButtonPuzzle ( RT ): The solver must hold a colored button for a specific number of seconds based on the button’s color, the strip’s color when pressed, and the time remaining on a timer.
  • ColorPuzzle ( MR , MSR ): The solver presses squares of the least common color in a 4x4 grid, then follows a sequence based on a table, aiming to turn all squares white.
  • KeypadPuzzle ( MG , MSR ): The solver must describe the symbol of each button in a 2x2 grid. The expert must then identify a column in the manual containing these four unique symbols and tell the solver to press the symbols in the correct order from top to bottom.
@@
  • MazePuzzle ( MG , MSR ): The solver navigates a mouse through a maze to a colored sphere, pressing the correct button to disarm the module based on the maze’s layout.
  • MemoryPuzzle ( MR , MSR ): The solver presses buttons according to specific positional and label-based rules over five stages, with incorrect presses resetting progress.
@@
  • DogPuzzle ( RT ): The solver is presented with an image containing 0-4 dogs. Based on the number of dogs in the image, the solver must press the submit button when the last digit of the timer matches the number of dogs in the image.
  • WhoPuzzle ( MG ): The solver must tell the value on a display to the expert, who will identify a button label. The solver must then tell this label to the expert, and then press the correct button based on a detailed list of instructions.
@@ 4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents or multimodal large language models (MLLMs). During our experiments, we limit the number of conversation turns to 20, allowing for a unified and systematic assessment of interactions. We use greedy decoding when available to maintain consistent agent output across runs and run inference on a single NVIDIA A100 GPU with 80GB RAM. We parse the solver’s chosen actions at each conversation turn using exact string matching, and use PyAutoGUI  (Sweigart, 2023) to directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
@@ 4.2 Evaluation Metrics
We meticulously recorded several key performance metrics through multiple iterations of the experiments described below:
  • Success Rate (SR): The solver agent is assigned a 0 or 1 value for each puzzle depending on if it completed it. These values are averaged across all puzzles to obtain the success rate.
  • Partial Success Rate (PSR): Because our benchmark includes puzzles with multi-step reasoning, some puzzles can have a more precise success rate evaluation. For these multi-step puzzles, we assign the solver a number between 0 and 100 to indicate its progress towards the solution, and average this number across puzzles to obtain partial success rate. For single-step puzzles, partial success rate is identical to success rate.
@@
  • Average Conversation Length (ACL): We count the number of conversation turns the Solver took to arrive at the solution, or default to the maximum of 20 if the solver failed. This count is averaged across all puzzles to get Average Conversation Length.
@@ 4.3 Models
  • Human: We conduct several experiments in which a human plays as the solver or expert to provide a strong baseline. As hiring participants was prohibitively expensive and time consuming, we role played as agents ourselves across 30 sampled puzzles as a preliminary study, and leave further human participation to future work.
  • InternVL  (Chen et al., 2024): A vision-language model by Shanghai AI Lab, designed for cross-modal tasks like visual question answering and image-text retrieval. We evaluate both the 26b and 8b variants of the model.
  • QwenVL  (Bai et al., 2023): We use version 2 of QWenVL (QWen-VL2), offering enhanced pretraining for improved performance on vision-language tasks. We use the 2b and 7b variants.
@@
  • GPT-4V: A version of OpenAI’s GPT-4, GPT-4V incorporates visual processing, enabling it to interpret both text and images.
  • GPT-4o: An optimized, faster, and more cost-effective variant of GPT-4, used for applications requiring speed and efficiency.
  • GPT-4o1: The most recent version of OpenAI’s GPT series models, which claims to have improved reasoning capability via internal chain of thought.

### v2

@@ Abstract
The rapid advances of multimodal agents built on large foundation models have largely overlooked their potential for language-based communication between agents in collaborative tasks. This oversight presents a critical gap in understanding their effectiveness in real-world deployments, particularly when communicating with humans. Existing agentic benchmarks fail to address key aspects of inter-agent communication and collaboration, particularly in scenarios where agents have unequal access to information and must work together to achieve tasks beyond the scope of individual capabilities. To fill this gap, we introduce a novel benchmark designed to evaluate the collaborative performance of multimodal multi-agent systems through language communication. Our benchmark features a variety of scenarios, providing a comprehensive evaluation across four key categories of agentic capability in a communicative collaboration setting. By testing both agent-agent and agent-human collaborations using open-source and closed-source models, our findings reveal surprising weaknesses in state-of-the-art models, including proprietary models like GPT-4o. Some of these models struggle to outperform even a simple random agent baseline in agent-agent collaboration and only surpass the random baseline when a human is involved.
Machine Learning, ICML
@@ 1 Introduction
The field of multimodal agents is experiencing rapid growth  (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024), with research efforts expanding at an unprecedented pace. However, amidst this growth, a critical gap in research has emerged: the lack of focus on collaborative work (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b) among multiple multimodal agents. Synergistic operation of such agents is a highly promising but largely unexplored domain. Language agents can collaboratively finish complex tasks such as software development  (Qian et al., 2024; Du et al., 2024) or even machine learning research  (Schmidgall et al., 2025) by assuming functional roles such as system designer, function generator, etc. Current research on multimodal agents (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024) has mainly focused on individual agent capabilities, neglecting the potential for inter-agent collaboration. This limitation is further compounded by existing benchmarks such as TheAgentCompany (Xu et al., 2024a), VisualWebArena (Koh et al., 2024) and MME-RealWorld (Zhang et al., 2024), which do not assess collaborative performance between agents. As a result, our ability to evaluate and improve multi-agent systems remains constrained, hindering progress in this crucial area.
@@
Several critical questions emerge in the context of multimodal agent collaboration. How can different agents effectively communicate multimodal information through language when they have varying levels of access to information? In scenarios where different agents possess diverse task-specific capabilities, how can they collaborate to accomplish objectives beyond the scope of any individual agent? These research settings remain largely uncharted and present significant challenges. Furthermore, the ability of agents to handle incomplete information is of paramount importance, particularly when working with sensitive data  (Li et al., 2024) (i.e. Agent application in healthcare where privacy concerns are critical  (Tang et al., 2024)). Exploration of these questions is crucial for advancing the field of multimodal agent collaboration. By addressing these challenges, we can expand the applicability of multimodal agents in real-world scenarios (Zhang et al., 2024), particularly those involving sensitive or restricted information.
@@
Figure 1: Overview of the interaction between the Solver and Expert agents in our benchmark. Both agents operate with structured input corresponding to working and episodic memory. The Solver receives an image of the puzzle state (working memory) and makes decisions based on the available actions described in the task prompt. The Expert, guided by instruction manuals (working memory), provides advice based on the Solver’s descriptions, such as indicating which buttons to press. The Solver can choose to execute actions by interacting with the environment or communicate with the Expert for further guidance. Their interaction is documented through a dialogue, showcasing the cooperation required to complete the task. Both agents engage in self-reflection by referencing the conversation history, which is continuously updated and incorporated into their input as episodic memory.
@@
Motivated by these aforementioned issues, we propose a novel benchmark for evaluating collaborative multimodal multi-agent frameworks to address critical gaps in current approaches (see  Figure 1). Our evaluation setting also simulates a scenario where an in-house agent with direct access to sensitive data (i.e., the AI solver) collaborates with external expert agents (i.e., the AI expert) to analyze information without compromising privacy. This evaluation setting simulates how we handle and extract insights from sensitive datasets across various domains.
@@
We assess multimodal multi-agent systems using a series of carefully designed collaborative puzzle games. Building on cognitive science research employing simple puzzles to assess cognitive ability  (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006), we design vision-language puzzles to evaluate the core cognitive abilities of multimodal agents, including learning, adaptation, and problem-solving. These scenarios typically involve two-player setups where agents have access to different, complementary information. (i.e., in a bomb defusal game, one agent possesses details about the bomb, while the other has access to a disarming manual). By employing such diverse and interactive scenarios, we aim to provide a thorough assessment of multimodal multi-agent performance.
@@
Our benchmark includes 10 distinct, easily customizable puzzles with thousands of unique solutions. We tested two different settings (AI-AI and AI-Human) and evaluated several popular multimodal models, including closed-source models (GPT-4V  (Achiam et al., 2023), GPT-4o  (Hurst et al., 2024)) and open-source models (QwenVL (Bai et al., 2023), and InternVL  (Chen et al., 2024)). Surprisingly, even powerful closed-source models such as the GPT series do not outperform a human solver, highlighting a potential growth area for future model development. Our contributions are as follows:
@@
  • Using COMMA, we carefully record conversations and performance metrics between state-of-the-art multimodal models such as QwenVL, InternVL, LLaMA, Gemini, GPT-4o, etc (Section 4).
  • We categorize the agent capabilities tested in our model and common failure modes, providing insight into potential future research directions for improving inter-agent communication (Section 5).
@@ 2 Related Work
There are many emergent agent collaboration works  (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b; Ghafarollahi & Buehler, 2024; Li et al., 2023; Wu et al., 2023) among multiple language agents. Multi-agent systems arise mainly in two different scenarios: (1) role-playing different task executors (e.g., software development requiring different roles of agents, such as program manager, software architect, programmer  (Du et al., 2024; Qian et al., 2024; Hong et al., 2024), scientific discovery simulation  (Wu et al., 2023), and social simulation  (Park et al., 2023; Gurcan, 2024; Park et al., 2024)); (2) communicating between agents with different pieces of information  (Wu et al., 2023; Li et al., 2023) (e.g., consulting experts without sharing some sensitive or confidential data. In our case, the AI solver has some private multimodal data, and the AI expert has domain-specific knowledge or instructions).
@@
Instruction-based agent benchmarks evaluate an agent’s capability of following a human instructions to finish a task (e.g., navigating on a website, interacting with an operating system  (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024)). However, our benchmark focuses more on a communication-based evaluation where two clients engage in multi-turn conversations to solve a task collaboratively.
@@ 3.1 Design Principles of the Benchmark
Our benchmark is inspired by the cooperative gameplay scenario in Keep Talking and Nobody Explodes (Games, 2015). In this game, two players work together to defuse a bomb under time pressure. One player, the defuser, can see the bomb but lacks the instructions to disarm it. The other player, the expert, has access to the bomb’s manual but cannot see the bomb itself. The players must rely on effective communication to exchange information, navigate challenges, and defuse the bomb.
We adapt this dynamic for our benchmark by shifting the focus to solving vision-language puzzles within a communication-based agent framework. To better reflect this broader scope, we rename the defuser role to “Solver,” emphasizing its general-purpose functionality across diverse tasks. As multimodal agent systems continue to gain traction, our benchmark aims to rigorously evaluate their reasoning, communication, and collaborative capabilities. Its design is grounded in the following core principles:
Agentic Architecture:
We carefully structure the Solver and Expert agents, drawing on the cognitive agent terminology from  Sumers et al. (2023). The Solver agent is equipped with working memory, which includes the task prompt, a screenshot of the puzzle, and direct feedback from the environment based on its actions. Additionally, the Solver has episodic memory in the form of conversation history, enabling it to learn from past mistakes and make informed decisions. The Expert agent follows a similar architecture but lacks access to environmental feedback or puzzle screenshots in its working or episodic memory. Instead, it relies solely on communication with the Solver to infer the puzzle state and provide guidance. Figure 1 provides an illustration of the interaction between the agents.
Intelligence Testing in Cognitive Science:
Our benchmark draws inspiration from the foundational principles of intelligence, often defined as the ability to learn from experience, adapt to the environment, and solve problems using cognitive skills (Kempf-Leonard, 2005). Cognitive science research has shown that even simple tests can effectively measure cognitive ability (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006). Standardized intelligence tests, such as MENSA (MENSA International, n.d.) and the Wechsler Intelligence Scale for Children (Wechsler, 1949), frequently employ simple puzzles to evaluate these skills. Building on this approach, our benchmark aims to assess the core cognitive capabilities of multimodal agents by creating simple vision-language puzzles tailored to test these abilities.
@@
A critical aspect of our benchmark is evaluating natural language communication between agents. Similar to how players in the original game exchange information verbally, agents in our framework must use language to share observations, clarify ambiguities, and reason about tasks. For the agents to succeed, they must display clarity, efficiency, and depth of communication, making it an essential factor in task completion.
Controllability:
Our framework is designed to allow for flexible difficulty and agent customizability by users. By granting agents access to user-defined functions and configuration files, it enables a modular environment for communication between multimodal agents, providing a robust platform to evaluate their intelligence. Future users can easily customize the framework by incorporating their own challenging puzzles and manuals, tailoring it to simulate more realistic and complex scenarios.
@@ 3.2 Categories of Agent Capability
Private Information (PR) Some puzzles challenge agents to withold information that might be sensitive and should not be shared through communication. This is a critical skill for embodied agents operating in real-world environments when dealing with proprietary data such as medical or personal financial records.
@@ 3.3 Tasks
  • ATMPuzzle ( PR ): The solver must navigate a bank interface and either make a withdrawal or deposit depending on the amount of their balance. The solver must not reveal private information such as their PIN number or balance amount while communicating with the expert.
  • TelehealthPuzzle ( PR ): The solver is in a health crisis situation and presented a private image of their skin and their background information (sourced from PAD-UFES-20  (Pacheco et al., 2020)). The solver must communicate with the expert to diagnose the skin disease and select the appropriate treatment plan, while taking care to not reveal any private health information.
  • ColorPuzzle ( MR , MSR ): The solver aims to turn all of the squares in a 4x4 grid white. At each step, the solver should press squares based on the frequencies of colors, following the rules specified in a table.
  • KeypadPuzzle ( MG , MSR ): The solver must describe the symbol of each button in a 2x2 grid. The expert must then identify a column in the manual containing these four unique symbols and tell the solver to press the symbols in the correct order.
@@
  • MazePuzzle ( MG , MSR ): The solver navigates a mouse through a maze to a colored sphere, pressing the correct button to disarm the module based on the layout.
  • MemoryPuzzle ( MR , MSR ): The solver presses buttons according to specific positional and label-based rules over five stages, with incorrect presses resetting progress. The rules for the current action depend on buttons pressed previously during the conversation.
@@
  • WhoPuzzle ( MG ): The solver must read out the value on a display to the expert, who will identify a button position to read from. The solver must then tell the expert the label of this button, and then press the correct button based on a detailed list of instructions.
@@ 4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. We use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle and run inference on a single NVIDIA A100 GPU with 80GB RAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
@@ 4.2 Evaluation Metrics
We recorded several key performance metrics through multiple iterations of the experiments described below:
  • Success Rate (SR): The solver agent is assigned a 0 or 100 value for each puzzle depending on the completion status. These values are averaged across all puzzles to obtain the success rate.
  • Partial Success Rate (PSR): Because our benchmark includes puzzles with multi-step reasoning, some puzzles can have a more precise success rate evaluation. For these multi-step puzzles, we assign the solver a number between 0 and 100 to indicate its progress towards the solution, and average this number across puzzles to obtain a partial success rate. For single-step puzzles, the partial success rate is identical to the success rate.
@@
  • Average Conversation Length (ACL): We count the number of conversation turns the Solver took to arrive at the solution, or default to the maximum of 10 if the solver failed. This count is averaged across all puzzles to get the Average Conversation Length.
@@ 4.3 Models
  • Human: We conduct experiments in which a human plays as the solver or expert to provide a strong baseline. As hiring participants was prohibitively expensive and time-consuming, we role-played as agents ourselves across 100 sampled puzzles as a preliminary study, and leave further human participation to future work.
  • InternVL  (Chen et al., 2024): A vision-language model designed for cross-modal tasks like visual question answering and image-text retrieval. We evaluate the 8b variant of the model.
  • QwenVL  (Bai et al., 2023): We use QwenVL-2, offering enhanced pretraining for improved performance on vision-language tasks. We use the 7b variant.
  • LLaMA 3.2  (Touvron et al., 2024): We use the 11b instruction-tuned version of LLaMA 3.2, the first LLaMA model to directly support multimodal input.
  • LLaVA  (Liu et al., 2024a): We use version 1.6 of LLaVA with a Mistral 7b language model backbone.
@@
  • GPT-4V  (Achiam et al., 2023): A version of OpenAI’s GPT-4, GPT-4V incorporates visual processing, enabling it to interpret both text and images.
  • GPT-4o  (Hurst et al., 2024): An optimized, faster, and more cost-effective variant of GPT-4, used for applications requiring speed and efficiency.
  • Gemini 2.0  (Anil et al., 2023): Google’s most recent LLM with a focus on agent capabilities supporting different input modalities such as vision, audio and text.